OpenStudy (anonymous):

how do you do the substitution method for 3x+2y=11 and x+y=4?

OpenStudy (lilg132):

y = 4 - x 3x + 2(4 - x) = 11 3x + 8 - 2x = 11 3x - 2x = 11 - 8 x = 3

OpenStudy (lilg132):

dont forget to click best response

OpenStudy (lilg132):

3(3) + 2y = 11 9 + 2y = 11 2y = 11 - 9 2y = 2 y = 1

OpenStudy (anonymous):

how do you find the substitution method for 5x-3y=5 and 10x+6y=20?

OpenStudy (lilg132):

you didnt click best response